### v3

@@ 4.2 Comparisons with other methods
PGCA (Datcu et al., 2023) Local (No mask) GR (ours) 56.7 71.4 88.1 61.0 68.0 74.7
SAR-KGPA Local GR (ours) 70.3 82.1 91.6 67.7 75.1 81.6
@@
PGCA (Datcu et al., 2023) Local (No mask) GR (ours) 76.4 80.4 83.8 77.4 81.6 84.2
SAR-KGPA Local GR (ours) 81.3 83.3 85.8 80.6 82.7 85.2
@@
PGCA (Datcu et al., 2023) Local (No mask) GR (ours) 51.8 60.0 73.7 52.7 58.3 66.1
SAR-KGPA Local GR (ours) 54.8 62.6 75.5 56.8 60.4 66.8
As shown in Table 3, we evaluated the SSL methods in few-shot classification tasks. The comparison methods are ImageNet, MAE (He et al., 2022), Feature Guided MAE (FG-MAE) (Wang et al., 2023c), I-JEPA (Assran et al., 2023), LoMaR (Chen et al., 2022a), PGCA (Datcu et al., 2023), and Our SAR-KGPA. Table 3 also gives the difference in the mask area and target features. The ImageNet migrated the supervised pre-training weights on the ImageNet as initialization. Following the PGCA using complex SAR image’s features as target signal, we used our SAR gradient features for pixel-level contrastive learning to test the effectiveness of SAR feature, i.e., mask image modeling is removed from SAR-KGPA.
Table 3 studies the result of different pre-trained models on fine-grained SAR target classification. Due to the large difference between SAR and RGB images, ImageNet weights are hard to obtain effect features with small samples with fine-tuning, and using linear probing can have a better result with an unsuitable initialization weight. MAE trained on the SAR dataset performs better than ImageNet with fine-tuning, but the speckle noise affects its representation quality. HOG is not an effective solution to the speckle noise because multiplicative noise leads to the virtual dots in the strong scattering region. It results in a performance degradation in FG-MAE with HOG features. Similarly, I-JEPA has the feature collapse, in which the learnable predictor fails to learn a good feature space under speckle noise. Therefore, we use gradient features with SAR domain knowledge as targets, and PGCA and SAR-KGPA have achieved a more significant performance than other methods. By combining mask image modeling and SAR domain knowledge, SAR-KGPA learns high-quality contextual features for low-quality noisy SAR data compared to PGCA.
Another interesting phenomenon is the range of contexts, where it can be found that LoMaR with local reconstruction can achieve better fine-tuning results in vehicle and ship compared to MAE. This result illustrates the need to consider the range of contextual information in remote-sensing images. However, due to a linear decoder, LoMaR is more susceptible to the effects of speckle noise in the pixels, leading to poorer linear probing results.
Although the proposed method achieves better results on the aircraft dataset, it is not as good as vehicle and ship results. It is because the aircraft reflects electromagnetic waves with its streamlined structure, resulting in SAR images that may lack the complete aircraft structure, with only some strong scattering points and shadow edges. Because of this, the generic representation in SAR images still needs to be further explored.
@@ 4.3 Visualization
Compared to supervised pre-training and contrastive learning, which have a local attention range in the high layer, ViT with MIM has various attention ranges in different layers (Xie et al., 2023b). However, Figure 4 shows that this characteristic of MIM is related to data property and pixel-level pretext task. We first notice that the deep layers of MAE mainly focus on global information due to large SAR image scenes, which differs from MIM’s modeling properties with ImageNet. Then, LoMaR with local reconstruction acquires the ability to model local information. This phenomenon illustrates that local information is easily lost in remote sensing images, and local reconstruction can better capture contextual details. In addition, PGCA, which removes the mask in SAR-KGPA and has a pixel-level contrast learning task, can acquire local information. It suggests that the property of modeling local and global information originates from the pixel-level pretext task. Finally, our SAR-KGPA with integrated MIM and PGCA can model local and global features. It can be found that the addition of MIM makes SAR-KPGA pay more attention to high-level global contextual information and more diminished attention to local edge features compared to PGCA. From the above discussions, the large remoting sensing image challenges for MIM to learn effective representation, which is why we chose the LoMaR method to solve this problem.
4.4 Ablation Study
Table 4: Results of different target feature for SAR-KGPA. The baseline is pixel value. LPF: low pass filter. HOG: histogram of oriented gradient. GR: gradient-by-ratio. The classification metric is average accuracy (%). Bold indicates the best result, underline indicates the next best result.
fine-tuning
Method MSTAR FUSAR-Ship SAR-ACD
Pixel 55.3 69.8 84.5 78.6 82.0 84.8 51.7 60.1 71.7
LPF (Liu et al., 2023) 57.7 72.7 86.5 78.4 81.9 84.9 51.9 61.2 72.8
HOG (Wei et al., 2022) 30.2 32.4 40.3 58.1 63.3 66.6 51.7 57.9 64.7
SAR-HOG (Song et al., 2016) 64.4 76.4 90.5 76.7 80.8 85.3 51.4 60.0 73.4
GR (Dellinger et al., 2014) 62.8 78.5 88.2 80.6 82.8 86.0 55.3 62.9 73.0
GR (ours) 70.3 82.1 91.6 81.3 83.3 85.8 54.8 62.6 75.5
linear probing
Method MSTAR FUSAR-Ship SAR-ACD
Pixel 57.7 64.4 72.2 78.0 81.0 84.0 54.4 58.9 64.4
LPF (Liu et al., 2023) 56.1 62.6 68.1 78.7 82.0 84.6 54.9 59.5 65.2
HOG (Wei et al., 2022) 33.3 37.8 41.6 52.6 58.3 61.5 49.2 53.1 56.7
SAR-HOG (Song et al., 2016) 58.9 68.1 77.1 79.3 81.8 85.0 54.6 59.3 66.5
GR (Dellinger et al., 2014) 62.8 72.3 77.6 82.0 82.9 85.7 54.7 60.9 66.6
GR (ours) 67.7 75.1 81.6 80.6 82.7 85.2 56.8 60.4 66.8
As for MIM based on SAR images, another key lies in extracting high-quality target features from low-quality and noisy SAR images. It needs to be combined with the SAR’s physical knowledge. Here, we analyze the target feature selection and the multi-scale setting.
Figure 5: Multi-scale kernel settings for our gradient-by-ratio. Here, the scale 1/2/3/4 is with $r$ equal to 5/7/13/17, and the multi-scale contains all four scales. Multi-scale is more suitable than single scale due to various targets in remote sensing images.
Figure 6: Scaling experiment of SAR-KGPA in dataset size, model parameters, and training epoch. We find that SAR-KGPA can benefit from these three aspects, but due to the limitation of dataset size, a large model and long training epochs lead to overfitting and reduce the linear probing performance.
Target feature selection. In Table 4, we compare the results of pixel value, Low Pass Filter (LPF) (Liu et al., 2023), SAR-HOG (Song et al., 2016), and Gradient-by-Ratio (GR) (Dellinger et al., 2014). All the changes are made only in the target features, while other settings remain unchanged.
The Pixel in Table 4 uses MAE’s Vit decoder, which suffers less noise interference than the LoMaR’s single projection head. The LPF can mitigate noise interference, but we observe a slight drop of LPF compared to Pixel on the MSTAR’s linear probing result. Therefore, inspired by the HOG feature in MaskFeat, we prefer to obtain high-quality target features further based on the suppressing noise.
Although HOG features have shown effectiveness in FG-MAE for SAR scene-level SSL, we found that the HOG features are unsuitable for target-level SSL in Table 4. This result is because speckle noise can make the gradient computation in small target regions appear to be many noise points, greatly affecting the quality of target features. Inspired by the physics-guided contrastive architecture using SAR physics features, we investigate the HOG feature variant in the SAR domain, i.e., SAR-HOG. The core of SAR-HOG is the change in the gradient calculation from differential gradient to ratio gradient. As shown in Table 4, we observe a significant improvement compared to HOG due to this modification. However, the sampling method of HOG may lead to the loss of some small target scattering points, such as the discrete points of an aircraft or a small boat. Therefore, we focus on the gradient information, i.e., target edges. Moreover, basic information, such as the shape and edges of an object, is a generic property (Liu et al., 2022; Cheng et al., 2022; Shi et al., 2020).
We consider two versions of GR in SAR. One is computing gradient with a linear kerenal (Song et al., 2016) whose value equals 1. Another is computing gradient with a Gaussian kernel (Dellinger et al., 2014) whose value obeys a Gaussian distribution. The Gaussian kernel can act as a filter, but the parameters need to be adjusted so as not to blur the target edges overly. For simplicity, we use a linear kernel without spending too much time tuning the parameters of the Gaussian kernel at different scales. Another important thing is that we propose a multi-scale setting instead of single-scale features due to the increased richness of targets included in the pre-training set.
Multi-scale setting. Multi-scale can improve the feature representation for various targets in remote sensing. Figure 5 gives the result of different kernel settings. Scale settings have various effects: smaller scales can better extract small target edges, while larger scales are more obvious for noise suppression. Hence, learning target features at different scales is more advantageous than a single scale.
@@ 4.5 Scaling Experiment
MIM can benefit from scaling data, model parameters, and computational resources, but it will suffer from overfitting when using small data and long training epochs on large models (Xie et al., 2023c). Therefore, we evaluate the scalability of SAR-KGPA with linear probing performance on SAR data. Figure 6 gives the scaling experiment in dataset size, model parameters, and training epoch.
@@
Training epoch. Similarly, too-long training epochs resulted in the overfitting of Vit-Base, and our default training setting is 200 epochs. Therefore, the next step is to collect more SAR target data to exploit the potential of SSL and foundation models.
@@ 5 Conclusion
In our work, we investigate the scale problem and target feature in MIM with SAR target images. By leveraging domain knowledge specific to SAR, we extend the applicability of MIM to SAR imagery, which is a type of noisy remote sensing data. Through extensive experimentation and analysis, we demonstrate the effectiveness of the proposed SAR-KGPA in obtaining high-quality feature representations for SAR ATR. Our research highlights the importance of embracing SSL to take advantage of the increasing availability of SAR data. The large pre-training samples give SSL the potential to achieve a foundation model for various downstream tasks across targets, scenes, and sensors.
However, it is worth noting that our pre-training dataset is not as extensive as ImageNet to fully exploit the potential of SSL for SAR ATR. In addition, it needs further investigation for other foundation model architectures, such as ConvNeXt. While our work primarily focuses on feature representation across different targets, it is essential to explore and discuss various tasks, including detection and segmentation. In the future, we plan to enhance our proposed scheme by improving data collection methods, refining model architectures, and exploring novel pre-training target features. Our ultimate goal is to successfully apply the foundation model with SSL in SAR target recognition.

### v4

@@ 4.3 Comparisons with other methods
SAR-JEPA Local $\rm{GR_{lin}}$ 70.3 82.1 91.6 67.7 75.1 81.6
@@
SAR-JEPA Local $\rm{GR_{lin}}$ 81.3 83.3 85.8 80.6 82.7 85.2
@@
SAR-JEPA Local $\rm{GR_{lin}}$ 54.8 62.6 75.5 56.8 60.4 66.8
Figure 5: Training loss and test accuracy curves for one fine-tuning of MSTAR 10-shot in Table 6. From the figure, effective self-supervised learning decreases the training loss rapidly and converges to a better-generalized region.
As shown in Table 6, we evaluated the SSL methods in few-shot classification tasks. The comparison methods are ImageNet, MAE (He et al., 2022), Feature Guided MAE (FG-MAE) (Wang et al., 2023d), I-JEPA (Assran et al., 2023), LoMaR (Chen et al., 2022a), and Our SAR-JEPA. Table 6 also gives the difference in the mask area and target features. The ImageNet migrated the supervised pre-training weights on the ImageNet as initialization.
Table 6 studies the result of different pre-trained models on fine-grained SAR target classification. Due to the large difference between SAR and RGB images, ImageNet weights are hard to obtain effect features with small samples under fine-tuning, and using linear probing can have a better result with an unsuitable initialization weight. MAE trained on the SAR dataset performs better than ImageNet with fine-tuning, but the speckle noise affects its representation quality. HOG is not an effective solution to the speckle noise because multiplicative noise leads to the virtual dots in the strong scattering region. It results in a performance degradation in FG-MAE with HOG features, and FG-MAE converges slower than MAE in Figure 5. Our method is a special form of the I-JEPA (Assran et al., 2023) method in SAR. But I-JEPA has the feature collapse, in which the learnable predictor fails to learn a good feature space under speckle noise. In Figure 5, I-JEPA converges to a region with poor test performance. Figure 5 also shows that effective self-supervised learning decreases the training loss rapidly and converges to a better-generalized region
Another interesting phenomenon is the range of contexts, where it can be found that LoMaR with local reconstruction can achieve better fine-tuning results in vehicle and ship compared to MAE. This result illustrates the need to consider the range of contextual information in remote-sensing images. However, due to a linear decoder, LoMaR is more susceptible to the effects of speckle noise in the pixels, leading to unstable performance in Table 6 and slow convergence rate in Figure 5.
Although the proposed method achieves better results on the aircraft dataset, it is not as good as vehicle and ship results. In particular, fine-tuning at SAR-ACD 10-shot is worse than the linear probing, indicating slight overfitting phenomena. It is because the aircraft reflects electromagnetic waves with its streamlined structure, resulting in SAR images that may lack the complete aircraft structure, with only some strong scattering points and shadow edges. Besides, our pre-training did not include many aircraft target samples. Because of this, the generic representation in SAR images still needs to be further explored.
@@ 4.4 Visualization
Compared to supervised pre-training and contrastive learning, which have a local attention range in the high layer, ViT with MIM has various attention ranges in different layers (Xie et al., 2023b). However, Figure 6 shows that this characteristic of MIM is related to data property and pixel-level pretext task. We first notice that the deep layers of MAE mainly focus on global information due to large SAR image scenes, which differs from MIM’s modeling properties with ImageNet. Then, LoMaR-SAR with local reconstruction acquires the ability to model local information. This phenomenon illustrates that local information is easily lost in remote sensing images, and local reconstruction can better capture contextual details. In addition, PGCA, which removes the mask in SAR-JEPA and has a pixel-level contrast learning task, can acquire local information. It suggests that the property of modeling local and global information originates from the pixel-level pretext task. Finally, our SAR-JEPA with integrated MIM and PGCA can model local and global features. It can be found that the addition of MIM makes SAR-JEPA pay more attention to high-level global contextual information and more diminished attention to local edge features compared to PGCA. Therefore, the differences in the properties between SAR and RGB images make the same self-supervised learning method show different properties, which is a great challenge for the SSL methods in SAR.
@@ 4.5 Scaling Experiment
Figure 7: Scaling experiment of SAR-JEPA in dataset size, model parameters, and training epoch. We find that SAR-JEPA can benefit from these three aspects, but due to the limitation of dataset size, a large model and long training epochs lead to overfitting and reduce the linear probing performance.
MIM can benefit from scaling data, model parameters, and computational resources, but it will suffer from overfitting when using small data and long training epochs on large models (Xie et al., 2023c). Therefore, we evaluate the scalability of SAR-JEPA with linear probing performance on SAR data. Figure 7 gives the scaling experiment in dataset size, model parameters, and training epoch.
@@
Training epoch. Similarly, too-long training epochs result in the overfitting of Vit-Base, and thereby, our default training setting is 200 epochs. Therefore, the next step is to collect more SAR target data to exploit the potential of SSL and foundation models.
@@ 5 Conclusion
In our work, we investigate the SSL for SAR ATR with our SAR-JEPA. SAR-JEPA combines two architectures to perform mask image modeling in gradient feature space for SAR images. By leveraging SAR domain features, we extend the applicability of MIM to SAR imagery, a type of noisy remote sensing data. Through extensive experimentation and analysis, we demonstrate the effectiveness of the proposed SAR-JEPA in obtaining high-quality representations for SAR ATR. Our research highlights the importance of embracing SSL to take advantage of the increasing availability of SAR data. The large pre-training samples. In the future, we will further extend the data and computational resources. give SSL the potential to achieve a foundation model for various downstream SAR ATR tasks across targets, scenes, and sensors.
However, it is worth noting that our pre-training dataset [^5] To the best of our knowledge, other SAR datasets can be further added to the pre-training, such as vehicle (Gotcha (Dungan et al., 2012), SIVED (Lin et al., 2023)), ship datasets (AIR-SARSHIP Sun et al. (2019), HRSID (Wei et al., 2020)), aircraft dataset (SADD Zhang et al. (2022)), and building datasets (FUSAR-Map Shi et al. (2021), OGSOD (Wang et al., 2023a)). However, how to further scale the SAR target data volume close to the million scale of ImageNet is a challenging problem. is not as extensive as ImageNet to fully exploit the potential of SSL for SAR ATR. In addition, it needs further investigation for other foundation model architectures, such as ConvNeXt. While our work primarily focuses on feature representation across different targets, it is essential to explore and discuss various tasks, including detection and segmentation. In the future, we plan to enhance our proposed scheme by improving data collection methods, refining model architectures, and exploring novel pre-training target features. Our ultimate goal is to successfully apply the foundation model with SSL in SAR target recognition.